Session: Sep 1999 (session 3)
Wavelength: 30 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (A. Kraus at Wed Sep 15 08:44:51 METDST 1999) Success
Formatter jumped 1 microsec at UT1154. Resync was done at UT1248.

Torun (K.Borkowski at Wed Sep 15 09:27:37 METDST 1999) Success
No known problems.

Jodrell (Alastair Gunn at Wed Sep 15 12:17:28 METDST 1999) Minor failures
Lovell telescope parked for essential engineering work at 07:45 UT. Did not restart observations, tape left running. Strong RFI. No other known problems.

Westerbork (Tony Foley at Thu Sep 16 10:56:58 METDST 1999) Minor failures
Telescope control computer crashed - losing a scan from 13:30 to 13:43 UT.

Calibration (Jean\-Francois Desmurs at Sun Nov 7 15:35:26 MET 1999) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: eb015b.antab. Read the comments in the file to get further information. (Rem : Tr no Tsys values, Check values of other stations.)